OpenStudy (anonymous):

what is 2+2 ? A.4 B.3 C.7 D.2

OpenStudy (anonymous):

7

OpenStudy (anonymous):

A. 4

OpenStudy (anonymous):

its 7 dont listen to them

OpenStudy (richyw):

don't listen to @Dianegrace. The answer is clearly 3

OpenStudy (anonymous):

yep its 7

OpenStudy (jlg030597):

A

OpenStudy (anonymous):

they are lying

OpenStudy (jlg030597):

This is insane a college student saying its 7

OpenStudy (anonymous):

comon bud stop liying its 7

OpenStudy (jlg030597):

lol haha its the easiest answer in the world its 7

OpenStudy (anonymous):

you guys arent very smart

OpenStudy (anonymous):

I see whats going on xD

OpenStudy (jlg030597):

Yup cause Obama says 4*3= 11

OpenStudy (anonymous):

XD omg..

OpenStudy (jlg030597):

Lol this is funny common the answers 3 we al know it is too

OpenStudy (jlg030597):

@xxraider365 am I right lol

OpenStudy (anonymous):

defintely 3

OpenStudy (anonymous):

i dont know bud im pretty sure its 7

OpenStudy (jlg030597):

Are you sure i got three and I did my work\[2+2=3\]

OpenStudy (anonymous):

I'm thinking it's 3 xD

OpenStudy (jlg030597):

See this is how I got it to Because the two variables are both 2 I figures that if a number repeats in a one step problem like this the second of the repeated numbers is subtracted by 1 lol hinse you get three.

OpenStudy (anonymous):

stupid obama

OpenStudy (jlg030597):

I hope that's a good way to do common core cause ill do it that way and then prove it just like this and as long as i can show how i got my answer i get full credit.

OpenStudy (richyw):

\[2+2 =\int_{-\infty }^{\infty } e^{-\frac{|x|}{2}} \, dx\]

OpenStudy (jlg030597):

Oh Lord i hope this aint a kid whos in Elementary school we'd have to explain what you just said @richyw

OpenStudy (anonymous):

im in college math and im a little confuse

OpenStudy (richyw):

i'm in college math and i'm a lot confused.

OpenStudy (anonymous):

lol